OpenStudy (anonymous):

A cylinder has a radius of 3 cm and a height of 8 cm. The surface area of the cylinder is _____ cm². 24pi 33pi 48pi 66pi

OpenStudy (anonymous):

The surface area of a cylinder is twice the area of the circular top: SA(top+bottom) = 2* pi*3^2 = 2*9pi = 18pi Added to the area of the rectangular sides of the cylinder, where the width is the height of the cylinder, and the length is the circumference of the circular top: SA(sides) = L*W = (2pi*r)*h = (2pi*3)*8 = 48pi So the surface area of the cylinder is: 18pi + 48pi = 66pi cm^2